OpenStudy (djanique):

differentiate y=x^4lnx-(1/2)x^2

OpenStudy (anonymous):

Use the chain rule to get \[y'=4x^3\ln x + x^4\tfrac{1}{x}-x\] which simplifies to \[y'=4x^3\ln x+x^3 - x\]

OpenStudy (anonymous):

y=x^4lnx-(1/2)x^2 y'=x^4/x+4x^3lnx-(1/2)2x y'=x^3+4x^3lnx-x